# Reproducing NevIR: Negation in Neural Information Retrieval

### v1

@@ Abstract.
Negation is a fundamental aspect of human communication, yet it remains a challenge for Language Models (LMs) in Information Retrieval (IR). Despite the heavy reliance of modern neural IR systems on LMs, little attention has been given to their handling of negation. In this study, we reproduce and extend the findings of NevIR, a benchmark study that revealed most IR models perform at or below the level of random ranking when dealing with negation. We replicate NevIR’s original experiments and evaluate newly developed state-of-the-art IR models. Our findings show that a recently emerging category—listwise Large Language Model (LLM) rerankers—outperforms other models but still underperforms human performance. Additionally, we leverage ExcluIR, a benchmark dataset designed for exclusionary queries with extensive negation, to assess the generalizability of negation understanding. Our findings suggest that fine-tuning on one dataset does not reliably improve performance on the other, indicating notable differences in their data distributions. Furthermore, we observe that only cross-encoders and listwise LLM rerankers achieve reasonable performance across both negation tasks.
@@
journalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
@@ 1. Introduction
Negation plays a crucial role in human communication, enabling the expression of contradictions, exclusions, and oppositions. Despite its common use in natural language, effectively handling negation remains a significant challenge for Language Models (LMs) (Ravichander et al., 2022). This issue has notable implications for Information Retrieval (IR) tasks, as different and complex forms of negation can appear in both user queries and documents.
This misinterpretation of negation is undesirable and can be harmful, particularly in high-risk retrieval tasks such as healthcare or legal settings, where it can lead to serious errors (Weller et al., 2024b). For example, as shown in Figure 1, misinterpreting negation in the medical question-answering domain can have severe and even catastrophic consequences.
@@
Despite its importance, the IR research community has paid little attention to addressing negation in retrieval. Existing work focuses on traditional term-based retrieval methods, overlooking more complex language structures such as negation (Averbuch et al., 2004; Kim et al., 2019). This is partly due to biases in existing collections, which lack sufficient examples of negation in the training data, leading to models’ inability to handle negation effectively.
@@
To address this data limitation, Weller et al. (2024b) proposes a novel negation benchmark – NevIR – to evaluate the performance of neural IR systems in handling negation. Their study tested models by ranking contrastive document pairs differing only in negation terms, using human-annotated queries to assess negation sensitivity. The results showed that most state-of-the-art (SotA) IR models struggle with negation, often performing worse than random ranking. Given the challenges identified in NevIR, an important question is whether its findings extend to newer families of IR models and LLM-based IR methods. More specifically, LLMs have demonstrated strong effectiveness in IR, making them promising candidates for retrieval tasks (Ma et al., 2024; Sun et al., 2023; Qin et al., 2024; Ma et al., 2023), but their ability to process negation remains uncertain.
A shortcoming of the original paper (Weller et al., 2024b) is that it relies on pairwise accuracy, a non-standard IR metric, limiting its broader applicability to IR tasks, such as ranking. In particular, the metric tests whether a model can differentiate negation between two given documents, offering insights into LMs’ handling of negation. However, the experiments in the original paper suggest that pairwise accuracy does not correlate with ranking-based metrics, such as nDCG@5. Therefore, the original study mainly addresses negation from an NLP perspective, focusing solely on language understanding, while missing the IR aspect of it — learning negation leads to diminished ranking performance. Another limitation of the paper is the absence of a baseline to assess the model’s overall performance on the data distribution. While the authors included a random baseline (25%) for the negation task, we argue that an additional baseline evaluating general performance on this distribution should have been provided. This would help distinguish whether the model’s poor performance is due to challenges in ranking documents with negation or an Out-Of-Distribution (OOD) effect.
In this work, we reproduce the NevIR paper (Weller et al., 2024b), as well as testing its generalisability by evaluating LLM-based retrieval methods. Ongoing research in NLP has also explored the understanding of LLMs’ negations, exploring factors such as model scaling and prompting techniques (Truong et al., 2023; Wei et al., 2022). By applying NevIR to this new class of models, we aim to understand whether existing conclusions about negation in IR hold or if recent advancements bring meaningful improvements. Our goal is to assess the extent to which negation remains a problem, given the latest advances in IR and NLP.
@@
Furthermore, Weller et al. (2024b) propose a preliminary solution for handling negation by fine-tuning IR models on the NevIR data set specific to negation. While fine-tuning has been shown to improve performance significantly, it remains unclear whether models learn negation patterns or overfit the data distribution of the NevIR dataset, particularly given that the dataset is created synthetically and could potentially lack negation and language diversity. To address this, we draw on the more recent work from Zhang et al. (2024) on ExcluIR, which includes a wider variety of negation types, and exclusionary queries. This comparison allows us to explore the generalisation of negation handling across different datasets and assess the effectiveness of NevIR in helping models understand negation in IR. Through this reproducibility study, we seek to extend the current work on negation in IR, contributing to a clearer understanding of how different datasets and models handle negation and ensuring that the insights gained are applicable beyond individual datasets. Furthermore, we aim to extend the initial NevIR experiments by exploring the potential trade-offs between improving negation understanding and preserving ranking performance.
By re-assessing the findings of the original work and extending them, we address the following research questions:
@@ 4. RQ4
Through our reproducibility work, we make the following findings:
  • We observe a positive trend in handling negation on newly developed SotA IR models. Notably, a newly emerged model category, listwise LLM re-rankers, achieved 20% gains compared to the previous approaches. We also observe that the recent advances in cross-encoders and bi-encoders with models such as GritLM and jina-re-ranker are also improving in negation understanding.
  • When fine-tuning, we find that, despite its superior base performance, LLM-based new architectures exhibit similar learning capabilities for the negation task as the more traditional models.
  • Fine-tuning experiments on NevIR and ExcliIR reveal that some models – cross-encoder – can generalise between the two negation datasets more effectively, going beyond the original data distribution.
@@
This research aims to assess the reproducibility of the original paper (Weller et al., 2024b) and the re-usability of the NevIR dataset to help negation understanding. We seek to verify its claims and assess its potential impact, contributing to the advancement of negation understanding in IR. Our code, which ensures full reproducibility of our experiments, is publicly available on GitHub [^1] https://github.com/thijmennijdam/NevIR-reproducibility.
@@ 2. Related Work
Negation in NLP. Negation is a complex semantic phenomenon in natural language that affects the performance of various Natural Language Processing (NLP) applications (Sineva et al., 2021; Morante and Blanco, 2021). It ”transforms an expression into another expression whose meaning is in some way opposed to the original” as mentioned by Morante and Blanco (2021). It occurs frequently, with the proportion of sentences with negation in English corpora ranging between 9% and 32% (Sineva et al., 2021). Negation is a challenging problem in machine translation, natural language inference, and sentiment analysis (García-Ferrero et al., 2023; Sineva et al., 2021).
@@
To highlight this problem, García-Ferrero et al. (2023) have created a dataset to assess LLMs handling of negation in NLP. They find that LLMs struggle with negative sentences and lack a deep understanding in negation. The importance of negation in NLP tasks is also highlighted by Hossain et al. (2022), which explores the role of negation in eight corpora for six popular natural language understanding tasks. The study finds that negation is virtually ignored by these corpora and that state-of-the-art transformers trained with these corpora face challenges in understanding negation.
Negation in Information Retrieval. At the time of the NevIR paper’s publication, little to no prior research had explored negation in neural IR models. More recently, Zhang et al. (2024) introduced ExcluIR, which includes an evaluation benchmark and a training set designed to improve retrieval models’ ability to handle exclusionary queries. While both NevIR and ExcluIR examine how IR models handle negation, they address different aspects of it: NevIR focuses on negation within documents, whereas ExcluIR emphasizes the exclusionary nature of queries. Nonetheless, it can be considered a branch of negation, rather than a completely separate class.
Additionally, Malaviya et al. (2023) developed QUEST, a dataset of natural language queries with implicit set operations, which map to a set of entities corresponding to Wikipedia documents, including negation. They analyzed several modern retrieval systems and, in alignment with the findings from NevIR, observed that IR models often struggle with queries involving negation
@@
LLMs for Information Retrieval. As LLMs have demonstrated strong representational capabilities, they are increasingly used as backbones in information retrieval (IR) architectures (Zhu et al., [n. d.]). These backbones are integrated into both bi-encoder and cross-encoder models, enhancing query-document representations for retrieval and reranking tasks (Zhu et al., [n. d.]). Around the introduction of NevIR (Weller et al., 2024b), most LLM-based IR models primarily relied on what are now considered relatively smaller-scale models, such as BERT (Kenton and Toutanova, 2019) and T5 (Raffel et al., 2020). With the emergence of larger and more accessible open-source LLMs, including those from LLaMA-3 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Qwen-2 (Yang et al., 2024), recent developments have integrated these models into IR systems, achieving state-of-the-art performance  (Weller et al., 2024a; Ma et al., 2024; Li et al., 2023; Muennighoff et al., 2024).
With these improvements in LLM performance, researchers have also begun using LLMs directly as rerankers, prompting them to reorder a list of retrieved documents given a query (Sun et al., 2023; Pradeep et al., 2023). RankGPT (Sun et al., 2023) has demonstrated that generative LLMs, when prompted effectively, can outperform traditional rerankers by leveraging their strong contextual understanding. Similarly, RankLLM (Pradeep et al., 2023) fine-tunes open-source LLMs for listwise document ranking, showing competitive performance with minimal additional training. To address the efficiency limitations of these models,  Gangi Reddy et al. (2024) introduce an approach that leverages the output logits of the first generated identifier to directly derive a ranked ordering of the candidates.
@@ 3. Overview of NevIR
NevIR was constructed by first identifying pairs of minimally different documents that include negation, using the CondaQA dataset (Ravichander et al., 2022) as a foundation. CondaQA is an English reading comprehension dataset designed to facilitate the development of models capable of processing negation effectively. It consists of 14,182 question-answer pairs with over 200 unique negation cues. The dataset was created through crowdworkers paraphrasing passages by rewriting negated statements in a way that preserves the sentence’s meaning.
Building on top of this, NevIR introduced a query generation step via Amazon Mechanical Turk. Crowdworkers were tasked with formulating one query per paragraph under specific constraints, ensuring that each query aligned most with one of the two minimally different documents.
@@ 4.1. Models
We evaluate nearly all the neural IR models [^3] Due to dependency issues with the RocketQA library, it was not possible to reproduce the RocketQA models (Qu et al., 2021) categorized and tested by Weller et al. (2024b), as well as update the benchmark with newly developed SotA models. Below, we describe each category of retrieval models and discuss the models evaluated within each. Additionally, we introduce listwise LLM re-ranking as a new model category, which emerged in recent work (Sun et al., 2023; Pradeep et al., 2023; Gangi Reddy et al., 2024).
@@
Late Interaction Models. Late interaction models encode queries and documents as vectors for each sub-word token. During inference, they compute similarity through a MaxSim operation. We tested ColBERTv1 (Khattab and Zaharia, 2020) and ColBERTv2 (Santhanam et al., 2022), as these remain representative of the state of the art in this category.
@@
Bi-Encoders. Bi-encoders encode queries and documents into single-vector representations, enabling fast similarity computations using dot products or cosine similarity. The original work evaluated diverse models, including some from SentenceTransformers (Reimers, 2019), DPR (Karpukhin et al., 2020) and CoCondenser (Gao and Callan, 2022). We extended this category by evaluating several additional models. DRAGON (Lin et al., 2023) is a model similar to DPR but adapted with advanced data augmentation techniques for enhanced generalisability. Promptriever (Weller et al., 2024a) introduces promptable retrieval and was released with both LLama-2-7B and Mistral-7B backbones. We also included OpenAI’s text-embedding-small and text-embedding-large models, and RepLlama (Ma et al., 2024), which uses a Llama-2 backbone. Furthermore, we evaluated the Qwen models with 1.5B and 7B variants (Yang et al., 2024), and GritLM (Muennighoff et al., 2024), which is based on a Mistral-7B backbone. GritLM is trained as a bi-encoder, cross-encoder, and generative model, but due to time constraints, we focused solely on its bi-encoder variant. Testing its cross-encoder and generative modes would be an interesting direction for future work.
@@
Cross-Encoders. Cross-encoders process both the document and the query together, offering highly expressive representations but at a higher computational cost. These models are typically used for re-ranking rather than retrieval. The original study evaluated various cross-encoders from SentenceTransformers (Reimers, 2019) and MonoT5 (Nogueira et al., 2020). In addition to these, we evaluated RankLlama (Ma et al., 2024), trained as a pointwise re-ranker with a Llama-2-7B backbone. We also tested various BGE models, including bge-reranker-base [^4] https://huggingface.co/BAAI/bge-reranker-base, bge-reranker-large [^5] https://huggingface.co/BAAI/bge-reranker-large, and bge-reranker-v2-m3666https://huggingface.co/BAAI/bge-reranker-v2-m3, which are highly par-ameter-efficient cross-encoders optimized for multilingual tasks. Finally, we included a Jina re-ranker, specifically jina-reranker-v2-base-multilingual [^7] https://huggingface.co/jinaai/jina-reranker-v2-base-multilingual, a re-ranking model designed for diverse cross-lingual retrieval tasks.
@@
Listwise LLM Re-rankers. In the Listwise LLM Re-ranking approach, LLMs, which are often pre-trained as foundation models, are adapted for ranking tasks. While these models are not explicitly fine-tuned for re-ranking in many cases, their rich contextual understanding and ability to assess semantic relevance enable them to reorder a given list of documents based on their alignment with a query. In this category, we evaluated RankGPT (Sun et al., 2023) and specifically tested GPT-4o-mini (OpenAI et al., 2024), GPT-4o (OpenAI et al., 2024), and o3-mini. Additionally, we included Qwen2 models with both 1.5B and 7B Instruct variants (Li et al., 2023), as well as Llama-3.2-3B-Instruct, Llama-3.1-7B-Instruct (Dubey et al., 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). These models span diverse parameter scales and are optimized for instruction-following tasks. For all LLMs, we employed standardized prompts from the RankGPT framework to ensure consistency across evaluations.

### v3

# by *[inlinelist,1]label=),itemjoin=, ,itemjoin*=, and

@@ Abstract.
Negation is a fundamental aspect of human communication, yet it remains a challenge for Language Models (LMs) in Information Retrieval (IR). Despite the heavy reliance of modern neural IR systems on LMs, little attention has been given to their handling of negation. In this study, we reproduce and extend the findings of NevIR, a benchmark study that revealed most IR models perform at or below the level of random ranking when dealing with negation. We replicate NevIR’s original experiments and evaluate newly developed state-of-the-art IR models. Our findings show that a recently emerging category—listwise Large Language Model (LLM) re-rankers—outperforms other models but still underperforms human performance. Additionally, we leverage ExcluIR, a benchmark dataset designed for exclusionary queries with extensive negation, to assess the generalisability of negation understanding. Our findings suggest that fine-tuning on one dataset does not reliably improve performance on the other, indicating notable differences in their data distributions. Furthermore, we observe that only cross-encoders and listwise LLM re-rankers achieve reasonable performance across both negation tasks.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730294isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrievalccs: Information systems Evaluation of retrieval resultsccs: Information systems Test collections
@@ 1. Introduction
Negation plays a crucial role in human communication, enabling the expression of contradictions, exclusions, and oppositions. Despite its common use in natural language, effectively handling negation remains a significant challenge for Language Models (LMs) (Ravichander et al., 2022). This issue has notable implications for Information Retrieval (IR) tasks, as different and complex forms of negation can appear in both user queries and documents. This misinterpretation of negation is undesirable and can be harmful, particularly in high-risk retrieval tasks such as healthcare or legal settings, where it can lead to serious errors (Weller et al., 2024b). For example, as shown in Figure 1, misinterpreting negation in the medical question-answering domain can have severe and even catastrophic consequences.
@@
Prior work on negation in IR, primarily from the pre-neural era, relied on rule-based and early statistical methods (Chapman et al., 2001; Harkema et al., 2009; Goldin and Chapman, 2003; Uzuner et al., 2011), focusing on explicit lexical cues. However, these methods largely overlooked more nuanced or implicit forms of negation. Despite its importance, there has been little research on how modern neural IR models handle negation. To address this gap, Weller et al. (2024b) propose a novel benchmark dataset —NevIR— to specifically evaluate how more recent neural IR systems handle negation. The NevIR dataset consists of pairs of queries, each semantically aligned with exactly one of two contrastive documents. These document pairs differ only in the presence or absence of negation terms (e.g., “The box is locked.” vs “The box is unlocked.”). The authors tested neural IR models by assessing their ability to correctly rank the appropriate document above its negated counterpart. The results showed that most state-of-the-art (SotA) IR models struggle with negation, often performing worse than random ranking. Given these limitations identified by the NevIR benchmark, an important question arises: do these findings generalise to newer IR model families, particularly LLM-based IR methods? Recently, LLMs have demonstrated strong effectiveness in IR, making them promising candidates for retrieval tasks (Ma et al., 2024; Sun et al., 2023; Qin et al., 2024; Ma et al., 2023), but their ability to process negation remains uncertain.
In this work, we therefore not only reproduce the experiments of the NevIR paper (Weller et al., 2024b), but also extend its evaluation to LLM-based retrieval methods. In doing so, we aim to assess whether recent advances in IR have improved negation handling or whether it remains a persistent challenge.
@@
Furthermore, Weller et al. (2024b) propose a preliminary solution for handling negation by fine-tuning IR models on the NevIR dataset. While this improves performance significantly, it remains unclear whether models truly learn negation patterns or overfit the data distribution of the NevIR dataset. This is particularly interesting given that the dataset is created synthetically and could potentially lack negation and language diversity. To address this, we build on the recent work of Zhang et al. (2024), who introduced ExcluIR—a benchmark focused specifically on exclusionary queries. For example, the query “What other flowers besides tulips grow in the Netherlands?” expresses exclusion by implicitly negating one category. By fine-tuning and evaluating models on both NevIR and ExcluIR, we explore the generalisation of negation handling across datasets and assess the effectiveness of NevIR in helping models understand negation in IR.
Through this reproducibility study, we seek to extend the current work on negation in IR, contributing to a clearer understanding of how different datasets and models affect negation understanding. Additionally, we aim to extend the initial NevIR experiments by exploring the potential trade-offs between improving negation understanding and preserving ranking performance. We address the following research questions:
@@ 4. RQ4
In response to these questions, we present the following findings:
  • We observe a positive trend in handling negation on newly developed SotA IR models. Notably, a newly emerged model category, listwise LLM re-rankers, achieved 20% gains compared to the previous approaches. We also observe that the recent advances in cross-encoders and bi-encoders with models such as GritLM and jina-reranker-v2-base-multilingual show an improvement in negation understanding.
  • When fine-tuning, we find that, despite its superior base performance, listwise LLM re-rankers have similar learning capabilities for the negation task as other model categories.
  • Fine-tuning experiments on NevIR and ExcluIR reveal that only the model from the cross-encoder category can generalise between the two negation datasets.
@@
Our code, which ensures full reproducibility of our experiments, is publicly available on GitHub [^1] https://github.com/thijmennijdam/NevIR-reproducibility.
@@ 2. Related Work
Negation in NLP. Negation is a complex semantic phenomenon in natural language that affects the performance of various Natural Language Processing (NLP) applications (Sineva et al., 2021; Morante and Blanco, 2021). It ”transforms an expression into another expression whose meaning is in some way opposed to the original” as mentioned by Morante and Blanco (2021). It occurs frequently, with the proportion of sentences with negation in English corpora ranging between 9% and 32% (Sineva et al., 2021). Negation is a challenging problem in machine translation, natural language inference, and sentiment analysis (García-Ferrero et al., 2023; Sineva et al., 2021). To highlight this problem, García-Ferrero et al. (2023) have created a dataset to assess LLMs handling of negation in NLP. They find that LLMs struggle with negative sentences and lack a deep understanding of negation. The importance of negation in NLP tasks is also highlighted by Hossain et al. (2022), which explores the role of negation in eight corpora for six popular natural language understanding tasks. The study finds that negation is virtually ignored by these corpora and that SotA transformers trained with these corpora face challenges in understanding negation.
Negation in Medical Information Retrieval. Negation is a well-recognised challenge in medical IR, as clinical notes frequently state the absence of conditions, which can mislead retrieval systems if not properly handled. Early rule-based tools like NegEx Chapman et al. (2001) and its extension ConText (Harkema et al., 2009) became standard components in clinical NLP pipelines, often integrated into IR systems such as those used in the TREC Medical Records track. These methods detect negation scopes around medical concepts and filter out negated mentions of query terms, significantly improving precision by reducing false positives. Subsequent machine-learning approaches—such as assertion classification in the i2b2 challenge Uzuner et al. (2011) and cohort retrieval by Taylor and Harabagiu Taylor and Harabagiu (2018)—further advanced negation handling. These methods, along with early neural models like NegBio Peng et al. (2018), perform reasonably well in narrow clinical settings but tend to plateau beyond simple cue patterns and require extensive domain-specific tuning.
Negation in Information Retrieval. Negation handling has also been explored in broader IR contexts, often through rule-based or hybrid methods. For instance, Goldin and Chapman (2003) demonstrated how negated terms can be filtered or reweighted in retrieval pipelines, showing that statistical models can effectively complement rule-based components. At the time of the NevIR paper’s publication, little research had tackled negation in neural IR models directly. More recently, however, Zhang et al. (2024) introduced ExcluIR, which provides an evaluation benchmark and training set aimed at improving how retrieval models handle exclusionary queries—negation expressed through user intent. While both NevIR and ExcluIR examine how IR models handle negation, they address different aspects of it: NevIR focuses on negation within documents, whereas ExcluIR emphasises the exclusionary nature of queries.Malaviya et al. (2023) developed QUEST, a dataset of natural language queries involving implicit set operations (including negation) that map to entities in Wikipedia. Their analysis aligns with findings from NevIR, showing that many modern retrieval models still struggle with negation-related phenomena.
@@
LLMs for Information Retrieval. As LLMs have demonstrated strong representational capabilities, they are increasingly used as backbones in information retrieval (IR) architectures (Zhu et al., [n. d.]). These backbones are integrated into both bi-encoder and cross-encoder models, enhancing query-document representations for retrieval and reranking tasks (Zhu et al., [n. d.]). Around the introduction of NevIR (Weller et al., 2024b), most LLM-based IR models primarily relied on what are now considered relatively smaller-scale models, such as BERT (Kenton and Toutanova, 2019) and T5 (Raffel et al., 2020). With the emergence of larger and more accessible open-source LLMs, including those from LLaMA-3 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Qwen-2 (Yang et al., 2024), recent developments have integrated these models into IR systems, achieving state-of-the-art performance  (Weller et al., 2024a; Ma et al., 2024; Li et al., 2023; Muennighoff et al., 2024). With these improvements in LLM performance, researchers have also begun using LLMs directly as rerankers, prompting them to reorder a list of retrieved documents given a query (Sun et al., 2023; Pradeep et al., 2023). RankGPT (Sun et al., 2023) has demonstrated that generative LLMs, when prompted effectively, can outperform traditional rerankers by leveraging their strong contextual understanding. Similarly, RankLLM (Pradeep et al., 2023) fine-tunes open-source LLMs for listwise document ranking, showing competitive performance with minimal additional training. To address the efficiency limitations of these models,  Gangi Reddy et al. (2024) introduce an approach that leverages the output logits of the first generated identifier to directly derive a ranked ordering of the candidates.
@@ 3. Overview of NevIR
NevIR was constructed by first identifying pairs of minimally different documents that include negation, using the CondaQA dataset (Ravichander et al., 2022) as a foundation. CondaQA is an English reading comprehension dataset designed to facilitate the development of models capable of processing negation effectively. It consists of 14,182 question-answer pairs with over 200 unique negation cues. The dataset was created through crowdworkers paraphrasing passages by rewriting negated statements in a way that preserves the sentence’s meaning. Building on top of this, NevIR introduced a query generation step via Amazon Mechanical Turk. Crowdworkers were tasked with formulating one query per paragraph under specific constraints, ensuring that each query aligned most with one of the two minimally different documents.
@@
Limitations. A shortcoming of the original NevIR paper was that it relies on pairwise accuracy, a non-standard IR metric, limiting its broader applicability to IR tasks, such as ranking. In particular, the metric tests whether a model can differentiate negation between two given documents, offering insights into LMs’ handling of negation. However, the experiments in the original paper suggest that pairwise accuracy does not correlate with ranking-based metrics, such as nDCG@5. Therefore, the original study mainly addressed negation from an NLP perspective, focusing solely on language understanding, while missing the IR aspect of it.
Another limitation of the paper was the absence of a baseline to assess the model’s overall performance on the data distribution. While the authors included a random baseline (25%) for the negation task, we argue that an additional baseline evaluating general performance on this distribution should have been provided. This would have helped distinguish whether the model’s poor performance is due to challenges in ranking documents with negation or an Out-Of-Distribution (OOD) effect.
@@ 4.1. Models
We evaluate nearly all the neural IR models [^3] Due to dependency issues with the RocketQA library, it was not possible to reproduce the RocketQA models (Qu et al., 2021) categorised and tested by Weller et al. (2024b), as well as update the benchmark with newly developed SotA models. Below, we describe each category of retrieval models and discuss the models evaluated within each. Additionally, we introduce listwise LLM re-ranking as a new model category, which emerged in recent work (Sun et al., 2023; Pradeep et al., 2023; Gangi Reddy et al., 2024).
@@
Late Interaction Models. Late interaction models encode queries and documents as vectors for each sub-word token. During inference, they compute similarity through a MaxSim operation. We reproducedColBERTv1 (Khattab and Zaharia, 2020) and ColBERTv2 (Santhanam et al., 2022), and did not test any new models as these remain representative of the state of the art in this category.
@@
Bi-Encoders. Bi-encoders encode queries and documents into single-vector representations, enabling fast similarity computations using dot products or cosine similarity. The original work evaluated a variety of models, including some from SentenceTransformers (Reimers, 2019), DPR (Karpukhin et al., 2020) and CoCondenser (Gao and Callan, 2022). We extended this category by evaluating several additional models. DRAGON (Lin et al., 2023) is a model similar to DPR but adapted with advanced data augmentation techniques for enhanced generalisability. Promptriever (Weller et al., 2024a) introduces promptable retrieval and was released with both LLama-2-7B and Mistral-7B backbones. We also included OpenAI’s text-embedding-small and text-embedding-large models, and RepLlama (Ma et al., 2024), which uses a Llama-2 backbone. Furthermore, we evaluated the Qwen models with 1.5B and 7B variants (Yang et al., 2024), and GritLM (Muennighoff et al., 2024), which is based on a Mistral-7B backbone. GritLM is trained as a bi-encoder, cross-encoder, and generative model, but due to time constraints, we focused solely on its bi-encoder variant. Testing its cross-encoder and generative modes would be an interesting direction for future work.
@@
Cross-Encoders. Cross-encoders process both the document and the query together, offering highly expressive representations but at a higher computational cost. These models are typically used for re-ranking rather than retrieval. The original study evaluated various cross-encoders from SentenceTransformers (Reimers, 2019) and MonoT5 (Nogueira et al., 2020). In addition to these, we evaluated RankLlama (Ma et al., 2024), trained as a pointwise re-ranker with a Llama-2-7B backbone. We also tested various BGE models, including bge-reranker-base [^4] https://huggingface.co/BAAI/bge-reranker-base, bge-reranker-large [^5] https://huggingface.co/BAAI/bge-reranker-large, and bge-reranker-v2-m3666https://huggingface.co/BAAI/bge-reranker-v2-m3, which are parameter-efficient cross-encoders optimised for multilingual tasks. Finally, we included a Jina re-ranker, specifically jina-reranker-v2-base-multilingual [^7] https://huggingface.co/jinaai/jina-reranker-v2-base-multilingual, a re-ranking model designed for diverse cross-lingual retrieval tasks.
@@
Listwise LLM Re-rankers. In the Listwise LLM Re-ranking approach, LLMs, which are often pre-trained as foundation models, are adapted for ranking tasks. While these models are not explicitly fine-tuned for re-ranking in many cases, their rich contextual understanding and ability to assess semantic relevance enable them to reorder a given list of documents based on their alignment with a query. In this category, we evaluated RankGPT (Sun et al., 2023) and specifically tested GPT-4o-mini (OpenAI et al., 2024), GPT-4o (OpenAI et al., 2024), and o3-mini. Additionally, we included Qwen2 models with both 1.5B and 7B Instruct variants (Li et al., 2023), as well as Llama-3.2-3B-Instruct, Llama-3.1-7B-Instruct (Dubey et al., 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). These models span diverse parameter scales and are optimised for instruction-following tasks. For all LLMs, we employed standardised prompts from the RankGPT framework to ensure consistency across evaluations.